Let X be a random variable with the following probability distribution:


x. -3 , 6 , 9


P(X=x). 1/6 , 1/2 , 1/3


Find ug(x), where g(X)= (2X+1)^2

1
Expert's answer
2022-03-28T16:33:08-0400

"\mu(g(X))=(-3*2+1)^2*\frac{1}{6}+(6*2+1)^2*\frac{1}{2}+(9*2+1)^2*\frac{1}{3}=209."